THE PECULIARITIES OF THE NMR SPIN-LATTICE RELAXATION IN PROTON EXCHANGED LINBO_{3}

\underline{IGOR VERTEGEL}, Institute of Physics NAS of Ukraine, Prospect Nauki 46, 03680 Kiev, Ukraine; EUGENY CHESNOKOV, Institute of Physics NAS of Ukraine, Prospect Nauki 46, 03680 Kiev, Ukraine; ALEXANDER OVCHARENKO, Institute of Physics NAS of Ukraine, Prospect Nauki 46, 03680 Kiev, Ukraine; and IVAN VERTEGEL, The Faculty of Mechanics and Mathematics, National Taras Shevchenko University of Kiev, Ukraine.

The temperature dependence of the spin-lattice relaxation time T_{1} of Li^{7} nuclei in the temperature range (170-430 K) was investigated in LiNbO_{3} polycrystalline samples: the clean and annealed ones in a hydrogen environment at temperature around 1000^{∘} C. The anomaly in the temperature dependence of T_{1} was found in range 300-340 K for both pure and annealed crystals. The reduction of the time T_{1} in the annealed lithium niobate crystal is caused by the creation of point defects (F^{+} or F-centers), with the dominant F-centers contribution. An increase in the activation energy in the annealed crystal can be explained by the following. It is known for the pure lithium niobate that an oscillation of lithium occurs in a symmetrical potential consisting of three wells. Formation of the oxygen vacancies in the annealed crystals is accompanied with extrinsic protons occupation of the vacancies. It leads to the symmetry violation and causes an i ncrease of the activation barrier.